1. The probability that it will be a sunny day is 0.4. Find the expected number of fine days in a week and also the standard deviation.


1
Expert's answer
2021-07-21T08:06:32-0400

"\mu=np=7*0.4=2.8."

"\sigma=\sqrt{np(1-p)}=\sqrt{7*0.4*0.6}=\sqrt{1.68}\approx1.2961."